A block is pulled to the right with a force of 250 N forming an angle of 42º, applying this force, the block reaches an acceleration of 1.5m / s.

The coefficient of friction is 0.3

a) Calculate the mass of the block.

b) Determine the normal force and the friction force


1
Expert's answer
2021-10-12T12:33:07-0400

a)

"F_{fr}=\mu N=\mu mg,"

"F\cos\alpha-F_{fr}=ma,"

"m=\frac{F\cos\alpha}{\mu g+a}=42~\text{kg},"

b)

"N=mg=412~\text N,"

"F_{fr}=\mu N=124~\text N."